OpenStudy (anonymous):

Find the general solution: (differential eqs - first order) ty'+2y=4t^2 y(1)=2

OpenStudy (anonymous):

i got the answer: y=(t^2)/3 + 5/6

OpenStudy (anonymous):

mu(t)= e^integral(2/t)dt

OpenStudy (anonymous):

your solution is incorrect. first divide whole equation with t y'+2/t y=4t find integrating factor \[\Large =e^{\int\limits_{}^{}\frac{2}{t}dt}=e^{2\ln(t)}=e^{\ln(t^2)}=t^2\] now Multiply whole equation with t^2 \[\Large t^2y'+2ty=4t^3\] can you do this now ?

OpenStudy (anonymous):

wouldnt it be t^2*y=integral(t^2 * 4t * dt)?

OpenStudy (anonymous):

nvm ur right

OpenStudy (anonymous):

do u mind giving me the answer so i can check at the end

OpenStudy (anonymous):

OS coc do not allow to give answers .. but we can work together to solve this ! let me know where you stuck. m here

OpenStudy (anonymous):

ok i integrated the eq u gave me and got t^3 * y + t^2 * y^2/2= t^4 + C

OpenStudy (anonymous):

(ty) + (y''t^2)?

OpenStudy (anonymous):

oops!!! the left side is product rule \[\Large t^2y'+2ty=4t^3\] can i write the left side as \[\Large \frac{d}{dt}(t^2y)=4t^3\] it it ok ?

OpenStudy (anonymous):

yes if you apply the product rule you will get the same y^2y'+2ty

OpenStudy (anonymous):

alright got it thx

OpenStudy (anonymous):

you are welcome :)